SpaceRef

SpaceRef


ISS On-Orbit Status 14 May 2002

Status Report From: NASA HQ
Posted: Tuesday, May 14, 2002

All ISS systems continue to function nominally, except as noted previously or below.  Another full work day aboard the station, 29 years after the launch of Skylab, first U.S. space station (1973).

Elektron update:  A commission to look into the failure of the Elektron oxygen generator is now at work in Moscow. Under investigation are two leading probable causes: (a) less likely: air in the system from an as yet unknown source, perhaps a CWC water container; (b) more likely: failure of the electronics box (signal and command matching unit, BSSK), which, if found faulty, could be replaced with an available spare on board.  MCC-M recommends continued use of the TGK solid-fuel oxygen generator (SFOG) candles for the next few days while troubleshooting goes on (according to MCC-M, there are no certified SFOG spares available on the ground at the present time). The Russians intend to launch the next two Progress cargo ships (8P, 9P) with a full oxygen load, each of which would last 12 days for a crew of three.  They are also looking into the possibility of carrying additional oxygen in the Progress' pressurized cargo bay (for which they apparently are not equipped at present), and they are investigating, jointly with MCC-H, other means of delivering stored O2 to the ISS.  If the current repair attempts are not successful, Moscow requests Shuttle return of the Liquid Unit, measuring 300 mm dia, 1100 mm high, 90 kg mass.

CDR Yuri Onufrienko again installed the Molniya-SM geophysics payload (GFI-10) at SM window 1, in time for Russian ground comm to initiate its operation.  Molniya-SM uses the twin-lens BFS-3M video-photometric system for the study of atmospheric, ionospheric and magnetospheric electromagnetic interaction related to storms and seismic activities.  Later today, after about ten hours of unattended operation, Onufrienko is to remove the equipment and stow it until the next session.

Onufrienko, assisted by FE-1 Carl Walz, then completed a major maintenance/repair task of removing and replacing one of the four SNT-50MP voltage and current stabilizer units (transformers) under the SM "floor".  To gain access to the system, they first had to remove the TVIS treadmill from the floor "pit", a one-hour task.  After replacing the failed SNT (A23), guided by ground specialists via S-band, Onufrienko connected the new unit to the BITS2-12 onboard telemetry system for checkout by the ground.  Finally, the TVIS was reinstalled in the "pit". [The four SNT-50MPs, at this stage of assembly, connect the Russian segment (RS) electrical power supply system (SZP) with the U.S. orbital segment (USOS), converting the 120 V dc electrical power coming from the PMA-1 (pressurized mating adapter #1) to 28.5 V dc used in the RS (after mission 12A, USOS voltage goes up to 160 V).]

Onufrienko, Walz and Bursch conducted a one-hour teleconference with the STS-111 and Expedition 5 crews, tagging up on the upcoming UF-2 mission, its timeline, getting EVA equipment and the Airlock (A/L) ready for the three spacewalks, including the A/L contingency tool kit contents and size, and the planned cargo transfers.  [After docking on Flight Day 3 (FD3), FEs Carl Walz and Sergey Treschev will officially "rotate" (swap places), followed on FD4 by the exchange of FE Dan Bursch with FE Peggy Whitson and CDR Yuri Onufrienko with CDR Valery Korzun.  MPLM Leonardo will be berthed on FD4; its unloading begins on FD5. The three EVAs, by Franklin Chang-Diaz and Philippe Perrin, will be conducted on FD5, FD7 and FD9. Their main objectives: install a PDGF (power and data grapple fixture) on the P6 module, transfer SM MMOD (micrometeoroid/orbital debris) shields from the Shuttle to a temporary location on PMA-1, install the MBS (mobile base system) on the MT (mobile transporter), and replace the wrist roll (WR) joint of the Canadian SSRMS with a new joint.]

Prior to the telecon, detailed instructions on EMU (extravehicular mobility unit) preparations for UF-2 were uplinked to the crew.  The U.S. spacesuits must be resized and fitted for their new wearers, which involves some rotation of suit hardware between suits.  The uplinked material was developed for the crew to keep track of this exacting swap-out.  Also uplinked were detailed pre-packing lists of EMU items and EVA tools for UF-2.

Walz and Bursch continued the final session of the renal (kidney) stone prevention research experiment, logging food and fluid intake and collecting urine samples during the day.  After each collection, the sample's ID was scanned with the bar code reader and the sample then stowed.

FE-2 Bursch continued his work on the BPS (biomass production system) test facility, today priming the Brassica root module for planting in PGC2 (plant growth chamber #2).  These are second-generation seeds from a seed pod, i.e., seeds from plants grown in microgravity, and it is of considerable scientific interest whether such seeds are capable of germination. [ Based on the Principal Investigator's observation of the Brassica plants from yesterdayâs video downlink, the seeds may be still be too young to germinate, but it is worth the effort because this opportunity will not be available post-flight.]

FE-1 Walz took radiation readings with the EVARM (EVA radiation monitoring) badges, as necessary to meet the EVA minus 21 days protocol requirement.

He also prepared the equipment for tomorrow's planned PHS (periodic health status) evaluation with blood labs. [This involves setting up the AMP (ambulatory medical pack) and the MEC (medical equipment computer) and opening the IFEP (in-flight examination program) on the MEC laptop.  The PHS is performed every 30 days by each crewmember and two weeks before landing and as clinically indicated.]

All crewmembers completed the weekly data collection for the NTXN "Interactions" interpersonal psychology program.

Other tasks completed were the regular daily SOSH life support system maintenance (Dan) and Inc. 4/8A payload status checks (Carl).

At 7:06 am EDT, Bursch conducted a ham radio session with Bordertown Primary School in Bordertown, South Australia, via telebridge stations in South Africa and Honolulu.  The school draws a large number of students from surrounding farms,- some students traveling up to 50 km one way. They had submitted 15 questions ahead of time ["How do you bathe in the Space Station?", and of course:  "Why do we need a Space Station?"]

A conjunction (close encounter) with an SL-8 rocket body (object #23279) is predicted for 5/16 (Thursday), with a TCA (time of closest approach) of 2:42 pm EDT. Currently estimated overall miss distance is 7 km; no avoidance maneuver is required at this time.  However, MCC-H and MCC-M ballistic specialists are going through initial planning for a maneuver to be ready, should it be necessary.

The U.S.-proposed test of using the CMGs with TA (control moment gyros with thruster assist) for maneuvering the ISS from the current flight attitude from earth-fixed LV/LH (local vertical/local horizontal) to the solar-oriented XPOP (x-axis perpendicular to orbit plane) remains scheduled for 5/17 (Friday).  MCC-Moscow (TsUP) concurs in the test which, for the first time, uses the CMGs instead of SM thrusters for the maneuver.

On 5/22 (Wednesday), TsUP plans to conduct an orbit correction/reboost maneuver, i.e. before UF-2, using Progress 7P. The burn, planned to yield a delta-V of 1m/s, was concurred in by Houston.

The U.S. CEO program had the following target areas today:  E. Mediterranean Dust and Smog (from the toe of the Italian boot ISS passed west of Greece and Crete and encountered the African coast at NW Egypt. A front had moved through the Mediterranean, clearing out the clouds, so the crew should have been able to look ahead and to the left for Saharan dust, which is still being carried out over the eastern Mediterranean), Lake Nasser, Toshka Lakes; Egypt (once past the Qattara Depression, crew was to look right to photograph the Toshka Lakes. The Aswan Dam was on track, then the pass went along the Red Sea coast over the coral reefs of the Dhalak Archipelago), W. Mediterranean Dust and Smog (conditions have cleared over France from Paris across the southern Rhone valley through the Maritime Alps. From the Riviera, crew was to look ahead [S] toward Corsica, then left along the Italian coast to record any newly developing aerosol masses over Italy. The volcanic island of Stromboli and Sicily, with Mt. Etna, were just to right of track), Rukwa Transform, Tanzania (another uncommonly clear period in south-central Africa  crew was asked to look right and take a mapping strip [10% overlap of frames] along the Lake Rukwa valley. Details of the Rungwe volcanic field were requested. Then, looking south down Lake Malawi to record the long, straight faults bounding the rift valley and its continuation into coastal Mozambique), Angolan Biomass Burning (fire season in Angola is under way, and weather is clear over the entire region. Crew to look left of track to record individual fires, where possible, as well as smoke palls), Industrialized Southeastern Africa (Johannesburg and Pretoria were just left of track; and the crew was to look for smog accumulations the valleys. The Vredefort meteor impact crater was just left of track. As ISS left the coast just south of Durban, they were to record any smog extending over the southern Indian Ocean), Salton Sea Water Levels (the crew should have had excellent weather as they reached Southern California, permitting them to take detailed photos of the Salton Sea. Their photos will be baseline data for significant changes in water availability that are expected there), Parana Wetlands, Post-Flooding (as ISS approached the great right-angle bend in the Parana River, the city of Rosario was just left of track. Of interest: looking left and documenting the condition of Parana wetlands following recent heavy rains. Also, taking near-nadir views of the upper estuary, where the Uruguay and Parana Rivers reach the estuary. Buenos Aires was just right of track and Montevideo, Uruguay just to the left).
CEO images can be viewed at the website http://eol.jsc.nasa.gov

U.S. and Russian Segment Status (as of 2:00 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is Off (failed); SFOG candles (TGKs) are being used.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA CO2 scrubber is Off.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) -- 759, temperature (deg C) -- 27.4, ppO2 (mmHg) -- 152.8, ppCO2 (mmHg) -- 2.5 (probably invalid)
SM Transfer Compartment:  Pressure (mmHg) -- 762, temperature (deg C) -- 20.5.
FGB Cabin:  Pressure (mmHg) -- 756, temperature (deg C) -- 20.3.
Node:  Pressure (mmHg) -- 750.25, temperature (deg C) -- 22.7 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab:  Pressure (mmHg) -- 752.49, temperature (deg C) -- 24.9, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 752.39, temperature (deg C) -- 22.4; shell heater temp (deg C) -- 22.6, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
PMA-1:  Shell heater temp (deg C) -- 23.0
PMA-2:  Shell heater temp (deg C) -- 17.7

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) have been removed for return to Earth.

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode.
SM batteries: Battery #6 is offline (failed); battery #7 is in cycle mode; all other batteries (6) are in "partial charge" mode.
FGB: Battery #2 is in cycle mode; all other batteries (5) are in "partial charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

Robotics:
SSRMS/Canadarm2 at Progress viewing position, with Keep Alive power on both strings.
RWS (robotics workstations) are Off.
SSRMS Prime string Wrist Roll (WR) joint to be replaced on UF-2.

ISS Orbit (as of this morning, 8:12 am EDT [= epoch]):
Mean altitude -- 388.3 km
Apogee -- 394.0 km
Perigee -- 382.7 km
Period -- 92.3 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0008328
Orbits per 24-hr. day -- 15.59
Altitude decrease -- 220 m (mean) in last 24 hours
Solar Beta Angle -- 18.4 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. '98) -- 19877
Current Flight Attitude -- LVLH (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7.25 deg., roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times, see
http://www.hq.nasa.gov/osf/station/viewing/issvis.htm

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.